For the compound FeCl3 , starting with 1.75 x 1024 chloride ions: please keep answers to two places past decimal point.


how many formula units would be possible to make?


how many moles of the compound would be present?


how many grams would be present ?


1
Expert's answer
2021-03-09T06:16:18-0500

No. Of molecules of FeCl3= (17.5×1023 )/ 3

= 5.83 ×1023

No. Of units 1 unit contains 6 molecule of FeCl3

=( 5.83 ×1023 )/ 6 = 0.97 ×1023


No of moles = (5.83 × 1023 )/ (6.022×1023)

= 0.96

Mass = 0.96 × 162.2 = 155.71g